Find the perimeter of a rectangular
region with side lengths 6 ft. and 12 ft. I am coming up with 36.
6 + 12 + 6 + 12
Am I figuring this correctly?

Answer by ankor@dixie-net.com(22740)

Yes but usually we write it:
p = 2(L) + 2(W)
p = 2(12) + 2(6)
p = 24 + 12
p = 36